Find the value of a for which

4cosec2  ( π (a + x)) +a2 - 4a = 0 has a real solution.

Grade:12

3 Answers

Priyansh Bajaj AskiitiansExpert-IITD
13 years ago

Dear Nirabhra,

Solution:- 4cosec2(n(a+x)) + a2 - 4a =0

[4cosec2(n(a+x)) -4] + (a2 - 4a + 4) =0

4[cosec2(n(a+x)) -1] + (a-2)2 =0

[2cot(n(a+x))]2 + (a-2)2 =0 which means a =2 and cot(n(a+x)) =0

Ans: a= 2

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

All the best Nirabhra!!!

Regards,
Askiitians Experts
Priyansh Bajaj

kirankumar kishor
13 years ago

 

Dear Nirabhra,

Solution:- 4cosec2(n(a+x)) + a2 - 4a =0

[4cosec2(n(a+x)) -4] + (a2 - 4a + 4) =0

4[cosec2(n(a+x)) -1] + (a-2)2 =0

[2cot(n(a+x))]2 + (a-2)2 =0 which means a =2 and cot(n(a+x)) =0

AA
5 years ago
 - 4a = 02 (π (a+x) ) + a24 cosec 

4 cosec 2 (π (a+x) ) =- a2 + 4a 

L .H .S is always greater than 4 so for real solution

 -a2 + 4a ≥4

or (a-2)2≤0

so only a=2 is possible